# PSDNorm: Test-Time Temporal Normalization for Deep Learning on EEG Signals

### v1

@@ Abstract
Distribution shift poses a significant challenge in machine learning, particularly in biomedical applications such as EEG signals collected across different subjects, institutions, and recording devices. While existing normalization layers, BatchNorm, LayerNorm and InstanceNorm, help address distribution shifts, they fail to capture the temporal dependencies inherent in temporal signals. In this paper, we propose PSDNorm, a layer that leverages Monge mapping and temporal context to normalize feature maps in deep learning models. Notably, the proposed method operates as a test-time domain adaptation technique, addressing distribution shifts without additional training. Evaluations on 10 sleep staging datasets using the U-Time model demonstrate that PSDNorm achieves state-of-the-art performance at test time on datasets not seen during training while being 4x more data-efficient than the best baseline. Additionally, PSDNorm provides a significant improvement in robustness, achieving markedly higher F1 scores for the 20% hardest subjects.
Machine Learning, ICML
@@ 1 Introduction
Machine learning techniques have achieved remarkable success in various domains, including computer vision, biology, audio processing, and language understanding. However, these methods face significant challenges when there are distribution shifts between training and evaluation datasets. For example, in biological data, such as electroencephalography (EEG) signals, the distribution of the data can vary significantly. Indeed, data is collected from different subjects, electrode positions, and recording conditions. This paper focuses on sleep staging, a task that consists of classifying the sleep stages of a subject based on EEG signals during one night of sleep. Depending on the dataset, the cohort can be composed of different age groups, sex repartition, health conditions, and recording conditions. Such variability brings shift in the distribution making challenging the generalization of the model to unseen datasets.
@@
Normalization layers are widely used in deep learning models to improve neural network training and generalization. The most commonly used normalization layers include BatchNorm (Ioffe & Szegedy, 2015), LayerNorm (Ba et al., 2016), and InstanceNorm (Ulyanov, 2016). BatchNorm computes statistics across the batch to normalize each feature, while LayerNorm normalizes all features within each individual sample. InstanceNorm applies normalization independently to each channel within a sample. Other normalization layers have been proposed for specific tasks such as (Kobler et al., 2022) for covariance matrices of EEG signals or (Kim et al., 2021) for time-series forecasting. However, these normalization layers have been shown to be insufficient to address data distribution shifts in the temporal correlation of signals. Hence, normalizations for sleep staging segmentation have been proposed as a pre-processing step either over an entire night of recordings (Apicella et al., 2023) or within smaller temporal windows (Chambon et al., 2018). However, recent studies (Gnassounou et al., 2023, 2024) show that taking into account temporal correlation is important when normalizing, and proposed the Temporal Monge Alignment (TMA) method, which aligns the Power Spectral Density (PSD) of signals to a common reference using Monge mapping.
@@
In sleep staging, numerous neural network architectures have been proposed to classify sleep stages. These models can process data in different formats, including raw signal inputs (Chambon et al., 2018; Perslev et al., 2021; Guillot & Thorey, 2021), spectrogram representations (Phan et al., 2023, 2019), or a combination of both (Phan et al., 2022). While many models are typically evaluated on a limited number of datasets, the work by (Perslev et al., 2021) introduced U-Sleep, a model trained on a large-scale dataset of sleep recordings. Their architecture, based on U-Time (Perslev et al., 2019), incorporates BatchNorm layers to mitigate data variability, employing a domain generalization approach: training a single model on a sufficiently diverse set of domains to ensure it generalizes to unseen datasets without additional adaptation. While this strategy shows promising performance, it raises the question of whether the model could be further adapted to the specifics of new subjects without retraining.
@@
Domain Adaptation methods have been proposed to address this issue and provide adaptation to new domains. Traditional DA methods typically align source and target data distributions in feature space during training (Sun & Saenko, 2016; Damodaran et al., 2018). In the context of sleep staging, DA has been used to address domain shifts (Eldele et al., 2021). However, this approach has limitations, as it requires access to the source domain for each new target DA. An alternative is test-time DA, where the model is adapted to the target domain during inference without requiring access to the source domain (Gnassounou et al., 2023). This approach offers greater flexibility, enabling clinical use of the model without retraining or sharing the source data.
@@
In this work, we introduce the PSDNorm deep learning layer, a novel approach to address distribution shifts in machine learning for signals. PSDNorm leverages Monge Mapping to incorporate temporal context and normalize feature maps effectively. As a test-time DA method, it adapts to distribution shifts during inference without extra training or source domain access. Compared to a normalization like LayerNorm or InstanceNorm, PSDNorm exploits the sequential nature of the intermediate layer representations as illustrated in Figure 1. We evaluate PSDNorm through extensive experiments on 10 sleep staging datasets. PSDNorm achieves state-of-the-art performance and requires 4 times fewer labeled data to match the accuracy of the best baseline. Moreover, it improves robustness with higher F1 scores for the 20% most challenging subjects, proving its effectiveness against domain shifts. Results highlight the potential of PSDNorm as a practical and efficient solution for tackling domain shifts in signals.
The paper is structured as follows: Section 2 discusses the U-Time model and existing normalization layers and pre-processing. Section 3 introduces PSDNorm, followed by numerical results in Section 4.
@@ 2 Related Works
In this section, we first present U-Time (Perslev et al., 2019, 2021), a convolutional deep learning model designed for temporal sequence segmentation. Then, we review fundamental concepts of normalization layers with the BatchNorm (Ioffe & Szegedy, 2015) and InstanceNorm (Ulyanov, 2016) layers. Finally, we recall the Temporal Monge Alignment (TMA) method (Gnassounou et al., 2023) that aligns the PSD of signals using optimal transport.
2.1 U-Time: CNN for time series segmentation
U-Time (Perslev et al., 2019, 2021) is a convolutional neural network (CNN) inspired by the U-Net architecture (Ronneberger et al., 2015), designed for segmenting temporal sequences. U-Time maps sequential inputs of arbitrary length to sequences of class labels on a freely chosen temporal scale. The architecture is composed of several encoder and decoder blocks, with skip connections between them.
Encoder blocks
Decoder blocks and Segmentation Head
The decoding part of U-Time is symmetrical to the encoding part. Each decoder block doubles the signal length and decreases the number of channels. It is composed of a convolutional layer, an activation function, a BatchNorm layer, an upsampling layer and a concatenation layer of the skip connection of the corresponding encoding block. Finally, the segmentation head applies two convolutional layers with an activation function in between to output the final segmentation. It should be noted that U-Time employs BatchNorm layers but other normalization layers, such as LayerNorm (Ba et al., 2016) or InstanceNorm (Ulyanov, 2016) are possible.
2.2 Normalization Layers
Over the last decade, numerous normalization layers have been proposed to improve the training and robustness of deep neural networks. Among them, the BatchNorm (Ioffe & Szegedy, 2015), InstanceNorm (Ulyanov, 2016), and LayerNorm (Ba et al., 2016) layers are the most popular. We present the first two as they are routinely used in deep learning models for time series processing, and illustrate them in Figure 1.
BatchNorm
The BatchNorm layer (Ioffe & Szegedy, 2015) normalizes features maps in a neural network to have zero mean and unit variance. At train time, given a batch =$\mathcal{B}=\{\mathbf{G}^{(1)},\dots,\mathbf{G}^{(N)}\}\subset\mathbb{R}^{c\times\ell}$ of $N$ pre-BatchNorm feature maps and for all $j,m,l\in\left\llbracket 1,N\right\rrbracket\,\times\left\llbracket 1,c\right\rrbracket\,\times\left\llbracket 1,\ell\right\rrbracket\,$, the BatchNorm layer is computed as
InstanceNorm and RevIN
Another popular normalization is the InstanceNorm layer (Ulyanov, 2016). During training, InstanceNorm operates similarly to (1), but the mean and variance are computed per sample instead of across the batch dimension, i.e., $\widehat{\mu}_{m}^{(j)}$ and $\widehat{\sigma}_{m}^{(j)}$ are computed for each sample $j$,
Hence, each sensor of each sample is normalized independently of the others. At test time, InstanceNorm behaves identically to its training phase and therefore does not rely on running statistics contrary to the BatchNorm. This layer has seen a growing interest for time-series forecasting with the RevIN layer (Kim et al., 2021) which applies InstanceNorm during the encoding and then reverses the normalization during the decoding to generate the forecast.
2.3 Power Spectral Density Alignment
Traditional normalization techniques, such as those discussed above, adjust statistical properties without accounting for the temporal structure of the data. To address this limitation, the Temporal Monge Alignment (TMA) (Gnassounou et al., 2023, 2024) was introduced as a pre-processing step to align temporal correlations by leveraging the Power Spectral density (PSD) of multivariate signals using Monge Optimal Transport mapping.
@@ 2.3 Power Spectral Density Alignment
Consider a multivariate signal $\mathbf{X}\triangleq[\mathbf{x}_{1},\dots,\mathbf{x}_{c}]^{\top}\in\mathbb{R}^{c\times\ell}$ of sufficient length. A standard assumption is that this signal follows a centered Gaussian distribution where sensors are uncorrelated and signals are periodic. This periodicity and uncorrelation structure implies that the signal’s covariance matrix is block diagonal, with each block having a circulant structure. A fundamental property of symmetric positive definite circulant matrices is their diagonalization (Gray, 2006) with real and positive eigenvalues in the Fourier basis $\mathbf{F}_{\ell}\in\mathbb{C}^{\ell\times\ell}$ of elements
@@
where $\mathbf{P}\in\mathbb{R}^{c\times\ell}$ contains positive entries corresponding to the PSD of each sensor. In practice, since we only have access to a single realization of the signal, the PSD is estimated with only $f\ll\ell$ frequencies, i.e., $\mathbf{P}\in\mathbb{R}^{c\times f}$. This amounts to considering the local correlation of the signal and neglecting the long-range correlations.
@@
The Welch estimator (Welch, 1967) computes the PSD of a signal by averaging the squared Fourier transform of overlapping segments of the signal. Hence, the realization of the signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ is decimated into overlapping segments $\{\mathbf{X}^{(1)},\dots,\mathbf{X}^{(L)}\}\subset\mathbb{R}^{c\times f}$ to estimate the PSD. The Welch estimator is defined as
@@
Let $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$ and $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(t)})$ be source and target centered Gaussian distributions respectively with covariance matrices following the structure (5) and PSDs denoted by $\mathbf{P}^{(s)}$ and $\mathbf{P}^{(t)}\in\mathbb{R}^{c\times f}$. Given a signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ such that $\operatorname{vec}(\mathbf{X})\sim\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$, the $f$-Monge mapping as defined by Gnassounou et al. (2023, 2024) is
$m_{f}\left(\mathbf{X},\mathbf{P}^{(t)}\right)\triangleq\mathbf{X}*\mathbf{H}\in\mathbb{R}^{c\times\ell}\;,$ (7)
where
$\mathbf{H}\triangleq\frac{1}{\sqrt{f}}\left({\mathbf{P}^{(t)}}\oslash{\mathbf{P}^{(s)}}\right)^{\odot\frac{1}{2}}\mathbf{F}_{f}^{*}\in\mathbb{R}^{c\times f}\;.$ (8)
@@
For Gaussian distributions admitting the decomposition (5), the Wasserstein barycenter (Agueh & Carlier, 2011) admits an elegant closed-form solution. Consider $K$ centered Gaussian distributions admitting the decomposition (5) of PSDs $\mathbf{P}^{(1)},\dots,\mathbf{P}^{(K)}$. Their barycenter is also a centered Gaussian distribution $\mathcal{N}(\mathbf{0},\overline{{\boldsymbol{\Sigma}}})$ admitting the decomposition (5) with PSD
@@ 3.1 Core Components of the layer
First, the estimation of the PSD of each feature map is performed using the Welch method. The per-channel mean $\widehat{{\boldsymbol{\mu}}}^{(j)}$ is computed for each feature map $\mathbf{G}^{(j)}$ as
$\widehat{{\boldsymbol{\mu}}}^{(j)}\triangleq\frac{1}{\ell}\sum_{l=1}^{\ell}\left[\mathbf{G}^{(j)}\right]_{:,l}\in\mathbb{R}^{c}\;.$ (10)
Then, the PSD of the centered feature map $\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top}$, denoted $\widehat{\mathbf{P}}^{(j)}$, is estimated as described in Equation (6). This centering step is required as feature maps are typically non-centered due to activation functions and convolution biases but they are assumed to have a stationary mean. The Welch estimation involves segmenting the centered feature map into overlapping windows, computing the Fourier transform of each window and then averaging them.
@@
Figure 2: Description of the running Riemanian barycenter. The barycenter of the batch $\widehat{\overline{\mathbf{P}}}_{\mathcal{B}}$ is estimate from the PSD of each batch sample. Then the running Riemanian barycenter is updated through an exponential average along the geodesic (- -), parameterized by $\alpha\in[0,1]$.
The PSDNorm aligns the PSD of each feature map to a barycenter PSD. This barycenter is computed during training by interpolating between the batch Wasserstein barycenter and the current running Riemanian barycenter using the geodesic associated with the Bures metric (Bhatia et al., 2019). The batch barycenter is first computed from the current batch PSDs $\big{\{}\widehat{\mathbf{P}}^{(1)},\dots,\widehat{\mathbf{P}}^{(N)}\big{\}}$ using Equation 9. To ensure gradual adaptation, the running barycenter is updated via an exponential geodesic average with $\alpha\in[0,1]$:
@@
$\widetilde{\mathbf{G}}^{(j)}$ $=m_{f}\left(\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top},{\widehat{\overline{\mathbf{P}}}}\right)$ (12)
$=\left(\left(\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top}\right)*\widehat{\mathbf{H}}^{(j)}\right)\in\mathbb{R}^{c\times\ell}$

### v2

# PSDNorm: Test-Time Temporal Normalization for Deep Learning in Sleep Staging

@@ Abstract
Distribution shift poses a significant challenge in machine learning, particularly in biomedical applications using data collected across different subjects, institutions, and recording devices, such as sleep data. While existing normalization layers, BatchNorm, LayerNorm and InstanceNorm, help mitigate distribution shifts, when applied over the time dimension they ignore the dependencies and auto-correlation inherent to the vector coefficients they normalize. In this paper, we propose PSDNorm that leverages Monge mapping and temporal context to normalize feature maps in deep learning models for signals. Notably, the proposed method operates as a test-time domain adaptation technique, addressing distribution shifts without additional training. Evaluations with architectures based on U-Net or transformer backbones trained on 10K subjects across 10 datasets, show that PSDNorm achieves state-of-the-art performance on unseen left-out datasets while being 4-times more data-efficient than BatchNorm.
@@ 1 Introduction
Machine learning techniques have achieved remarkable success in various domains, including computer vision, biology, audio processing, and language understanding. However, these methods face significant challenges when there are distribution shifts between training and evaluation datasets [1]. For example, in biological data, such as electroencephalography (EEG) signals, the distribution of the data can vary significantly. Indeed, data is collected from different subjects, electrode positions, and recording conditions. This paper focuses on sleep staging, a clinical task that consists in classifying periods of sleep in different stages based on EEG signals [2]. Depending on the dataset, the cohort can be composed of different age groups, sex repartition, health conditions, and recording conditions [3, 4, 5]. Such variability brings shift in the distribution making it challenging for the model to generalize to unseen datasets.
@@
Normalization layers are widely used in deep learning to improve training stability and generalization. Common layers include BatchNorm [6], LayerNorm [7], and InstanceNorm [8], which respectively compute statistics across the batch, normalize across all features within each sample, and normalize each channel independently within a sample. Some normalization methods target specific tasks, such as EEG covariance matrices [9] or time-series forecasting [10], but they do not fully address spectral distribution shifts reflected in the temporal auto-correlations of signals. In sleep staging, a simple normalization is often applied as preprocessing, e.g., standardizing signals over entire nights [11] or short temporal windows [12]. Recent studies [13, 14] highlight the importance of considering temporal correlation and spectral content in normalization, proposing Temporal Monge Alignment (TMA), which aligns Power Spectral Density (PSD) to a common reference using Monge mapping, going beyond simple z-score normalization. However, these methods remain preprocessing steps that cannot be inserted as layers in the network architecture as it is done with BatchNorm, LayerNorm or InstanceNorm.
@@
Sleep staging has been addressed by various neural network architectures, which process raw signals [12, 15, 16], spectrograms [17, 18], or both [19]. More recent approaches involve transformer-based models that handle multimodal [20], spectrogram [21], or heterogeneous inputs [22], offering improved modeling of temporal dependencies. However, most existing models are trained on relatively small cohorts, typically consisting of only a few hundred subjects, which limits their ability to generalize to diverse clinical settings. Notable exceptions include U-Sleep [15], which was trained on a large-scale dataset and incorporates BatchNorm layers to mitigate data variability, and foundational models [23, 24, 25] that achieve strong generalization from vast amount of data but require significant computational resources and are challenging to adapt without fine-tuning. Our focus is on developing smaller, efficient models that balance good generalization with ease of training and deployment in clinical practice.
@@
Domain generalization involves training models on data from multiple domains without using domain labels, with the goal of achieving broad generalization. In contrast, domain adaptation (DA) utilizes domain information to improve performance by aligning source and target distributions during training [26, 27]. While DA methods have been applied to address domain shifts in sleep staging [28], they typically require access to source data for each new target, which can be limiting. Test-time DA offers a more flexible approach, adapting the model during inference without requiring access to source data [13]. This makes test-time DA particularly well-suited for clinical applications where retraining or data sharing is impractical.
@@
In this work, we introduce the PSDNorm deep learning layer, a novel approach to address distribution shifts in machine learning for signals. PSDNorm leverages Monge Mapping to incorporate temporal context and normalize feature maps effectively. As a test-time DA method, it adapts to distribution shifts during inference without extra training or source domain access. Compared to a normalization like LayerNorm or InstanceNorm, PSDNorm exploits the sequential nature of the intermediate layer representations as illustrated in Figure 1. We evaluate PSDNorm through extensive experiments on 10 sleep datasets. This evaluation covers 10M of samples across 10K subjects, using a leave-one-dataset-out (LODO) protocol with 3 different random seeds. To the best of our knowledge, such a large-scale and systematic evaluation has never been conducted before. PSDNorm achieves state-of-the-art performance and requires 4 times fewer labeled data to match the accuracy of the best baseline. Results highlight the potential of PSDNorm as a practical and efficient solution for tackling domain shifts in signals.
The paper is structured as follows: Section 2 discusses existing normalization layers and pre-processing. Section 3 introduces PSDNorm, followed by numerical results in Section 4.
@@ 2 Related Works
In this section, we first review fundamental concepts of normalization layers. Then, we recall the Temporal Monge Alignment (TMA) method [13] that aligns the PSD of signals using optimal transport.
Normalization Layers
Normalization layers enhance training and robustness in deep neural networks. The most common are BatchNorm [6], InstanceNorm [8], and LayerNorm [7]. BatchNorm normalizes feature maps using batch and time statistics, ensuring zero mean and unit variance. The output is adjusted with learnable parameters. InstanceNorm normalizes each channel per sample using its own statistics, independent of the batch (see Fig. 1). Popular in time-series forecasting, it is used in RevIN [10], which reverses normalization after decoding. LayerNorm normalizes across all channels and time steps within each sample, with learnable scaling and shifting. While these normalization layers are widely employed, they operate on vectors ignoring statistical dependence and autocorrelation between their coefficients, which are prevalent when operating on time-series. To address this limitation, the Temporal Monge Alignment (TMA) [13, 14] was introduced as a pre-processing step to align temporal correlations by leveraging the Power Spectral density (PSD) of multivariate signals using Monge Optimal Transport mapping.
@@
Consider a multivariate signal $\mathbf{X}\triangleq[\mathbf{x}_{1},\dots,\mathbf{x}_{c}]^{\top}\in\mathbb{R}^{c\times\ell}$ of sufficient length. A standard assumption is that this signal follows a centered Gaussian distribution where sensors are uncorrelated and signals are periodic. This periodicity and uncorrelation structure implies that the signal’s covariance matrix is block diagonal, with each block having a circulant structure. A fundamental property of symmetric positive definite circulant matrices is their diagonalization [29] with real and positive eigenvalues in the Fourier basis $\mathbf{F}_{\ell}\in\mathbb{C}^{\ell\times\ell}$ of elements
@@
where $\mathbf{P}\in\mathbb{R}^{c\times\ell}$ contains positive entries corresponding to the Power Spectral Density of each sensor. In practice, since we only have access to a single realization of the signal, the PSD is estimated with only $F\ll\ell$ frequencies, i.e., $\mathbf{P}\in\mathbb{R}^{c\times F}$. This amounts to considering the local correlation of the signal and neglecting the long-range correlations.
@@
The Welch estimator [30] computes the PSD of a signal by averaging the squared Fourier transform of overlapping segments of the signal. Hence, the realization of the signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ is decimated into overlapping segments $\{\mathbf{X}^{(1)},\dots,\mathbf{X}^{(L)}\}\subset\mathbb{R}^{c\times F}$ to estimate the PSD. The Welch estimator is defined as
@@
Let $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$ and $\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(t)})$ be source and target centered Gaussian distributions respectively with covariance matrices following the structure (2) and PSDs denoted by $\mathbf{P}^{(s)}$ and $\mathbf{P}^{(t)}\in\mathbb{R}^{c\times F}$. Given a signal $\mathbf{X}\in\mathbb{R}^{c\times\ell}$ such that $\operatorname{vec}(\mathbf{X})\sim\mathcal{N}(\mathbf{0},{\boldsymbol{\Sigma}}^{(s)})$, the $F$-Monge mapping as defined by [13, 14] is
$m_{F}\left(\mathbf{X},\mathbf{P}^{(t)}\right)\triangleq\mathbf{X}*\mathbf{H}\in\mathbb{R}^{c\times\ell},\quad\text{where}\quad\mathbf{H}\triangleq\frac{1}{\sqrt{F}}\left({\mathbf{P}^{(t)}}\oslash{\mathbf{P}^{(s)}}\right)^{\odot\frac{1}{2}}\mathbf{F}_{F}^{*}\in\mathbb{R}^{c\times F}\;.$ (4)
@@
For Gaussian distributions admitting the decomposition (2), the Wasserstein barycenter [31] admits an elegant closed-form solution. Consider $K$ centered Gaussian distributions admitting the decomposition (2) of PSDs $\mathbf{P}^{(1)},\dots,\mathbf{P}^{(K)}$. Their barycenter is also a centered Gaussian distribution $\mathcal{N}(\mathbf{0},\overline{{\boldsymbol{\Sigma}}})$ admitting the decomposition (2) with PSD
@@ 3.1 Core Components of the layer
First, the estimation of the PSD of each feature map is performed using the Welch method. The per-channel mean $\widehat{{\boldsymbol{\mu}}}^{(j)}$ is computed for each feature map $\mathbf{G}^{(j)}$ as $\widehat{{\boldsymbol{\mu}}}^{(j)}\triangleq\frac{1}{\ell}\sum_{l=1}^{\ell}\left[\mathbf{G}^{(j)}\right]_{:,l}\in\mathbb{R}^{c}\;.$ Then, the PSD of the centered feature map $\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top}$, denoted $\widehat{\mathbf{P}}^{(j)}$, is estimated as described in Equation (3). This centering step is required as feature maps are typically non-centered due to activation functions and convolution biases but they are assumed to have a stationary mean. The Welch estimation involves segmenting the centered feature map into overlapping windows, computing the Fourier transform of each window and then averaging them.
@@
Figure 2: Description of the running Riemanian barycenter. The barycenter of the batch $\widehat{\overline{\mathbf{P}}}_{\mathcal{B}}$ is estimated from the PSD of each batch sample. Then the running Riemanian barycenter is updated through an exponential average along the geodesic (- -), parameterized by $\alpha\in[0,1]$.
The PSDNorm aligns the PSD of each feature map to a barycenter PSD. This barycenter is computed during training by interpolating between the batch Wasserstein barycenter and the current running Riemanian barycenter using the geodesic associated with the Bures metric [32]. The batch barycenter is first computed from the current batch PSDs $\big{\{}\widehat{\mathbf{P}}^{(1)},\dots,\widehat{\mathbf{P}}^{(N)}\big{\}}$ using Equation 5. To ensure gradual adaptation, the running barycenter is updated via an exponential geodesic average with $\alpha\in[0,1]$:
@@
$\widetilde{\mathbf{G}}^{(j)}$ $=m_{F}\left(\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top},{\widehat{\overline{\mathbf{P}}}}\right)=\left(\left(\mathbf{G}^{(j)}-\widehat{{\boldsymbol{\mu}}}^{(j)}\mathbf{1}_{\ell}^{\top}\right)*\widehat{\mathbf{H}}^{(j)}\right)\in\mathbb{R}^{c\times\ell}$ (7)